# Nonequilibrium thermodynamics of populations of weakly-coupled low-temperature-differential Stirling engines with synchronous and asynchronous transitions

### v1

@@ III.3 Justification of the conceptual model
In this section, we show that the differential equations describing the motion of rotating bodies and the heat flux equation of the conceptual model constructed in subsection B can be obtained by eliminating fast oscillating variables in those of the original coupled Stirling engine model. Since conditions (5) and (7) ensure that the time scales of $\omega_{i}$ and $\theta_{i}-\theta_{j}$ are sufficiently large compared to that of $\theta_{i}$, i.e., $\omega_{i}$ and $\theta_{i}-\theta_{j}$ change only slightly during which $\theta_{i}$ varies by $2\pi$, if we replace the time derivative of $\omega_{i}$ by its time average during a one-round change in $\theta_{i}$, $\omega_{i}$ and $\theta_{i}-\theta_{j}$ can be regarded as constants so that we can just replace the driving force portion of the engine by its time average over one cycle of $\theta_{i}$ in Eq. (3b):
@@
The higher-order terms produced by the averaging approximation of the heat flux can be ignored given that conditions (5) and (7) are satisfied. The conceptual model is thus expected to be reasonable as giving a physical interpretation given that conditions (5), (7) and (8) are satisfied.
The above discussion confirms the validity of the conceptual model constructed in subsection B, but there are still several issues to be addressed. One is that the results of the analysis in subsection B using the conceptual model do not agree with the results of the analysis in subsection A using the quasilinear response relations: the averaged power $\Bigl{\langle}P_{\mathrm{load}}^{\mathrm{D}}\Bigr{\rangle}$ applied by the flywheels ${\mathrm{D}}_{\mathrm{F}W}^{(1)},{\mathrm{D}}_{\mathrm{F}W}^{(2)},\cdots,{\mathrm{D}}_{\mathrm{F}W}^{(N)}$ to the loads (or the averaged power $-\Bigl{\langle}P_{\mathrm{fric}}^{\mathrm{D}}\Bigr{\rangle}$ due to the viscous friction) is given by Eq. (44) with the sign reversed:
@@
which corresponds to the averaged brake power $\langle P_{\mathrm{rel}}\rangle$ due to the relative motion of the weakly-coupled Stirling engines, but this result is different from the approximation $\langle P_{\mathrm{rel}}\rangle\approx-L_{11}^{-1}\left(\sum_{i=1}^{N}\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}^{2}\right)$ given by Eq. (27). The reason why the analysis based on the conceptual model has different results from the analysis using the quasilinear response relations will be discussed in subsection E. Another issue is the difference in energy dissipation due to viscous friction between the conceptual model and the original coupled engine model: the averaged energy dissipation rate due to viscous friction of the coupled Stirling engine is given by $\sum_{i=1}^{N}\Gamma\omega_{i}^{2}$, whereas in the conceptual model it is given by $\left(\Gamma+\frac{\sigma^{2}}{G}\Biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\Biggr{\rangle}_{\theta}\right)\left(\sum_{i=1}^{N}\omega_{i}^{2}\right)$. In Appendix B, we will see that the averaged energy dissipation rate $\frac{\sigma^{2}}{G}\Biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\Biggr{\rangle}_{\theta}\left(\sum_{i=1}^{N}\omega_{i}^{2}\right)$ in the conceptual model comes from the heat leakage of the coupled Stirling engines.
@@ III.4 Numerical experiments
Figure 5: Dependence of the coupling strength $K$ on the averaged brake power due to relative motion $\langle P_{\mathrm{rel}}\rangle$, calculated for coupled system models (3a)-(3b). The true values (true—) were calculated directly from the real data of model (3a)-(3b), while the approximate values (approx1—, approx2— and approx3—) were calculated from Eqs. (24), (26), and (27), respectively. The values of $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ and $\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$ contained in $\bigl{\langle}P_{\mathrm{rel}}\rangle$ were obtained from numerical experiments in the calculation of approximate values. (a): forward process, (b): backward process.
@@
Figure 6: Dependence of the coupling strength $K$ on the averaged brake power $\langle P_{\mathrm{load}}\rangle$, calculated for coupled system models (3a)-(3b). The true values (true—) were calculated directly from the real data of model (3a)-(3b). The values of $\langle\omega_{\mathrm{m}}\rangle$ required for the calculation of the approximate values were computed using analytical solutions obtained from the quasilinear response relations (17), while $\langle P_{\mathrm{rel}}\rangle$ was calculated in three different ways, using Eqs. (24) (corresponding to approx1—), (26) (corresponding to approx2—) and (27) (corresponding to approx3—). The values of $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ and $\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$ contained in $\bigl{\langle}P_{\mathrm{rel}}\rangle$ were obtained from numerical experiments in the calculation of approximate values. (a): forward process, (b): backward process.
@@
Figure 7: Dependence of the coupling strength $K$ on the thermal efficiency $\eta$, calculated for coupled system models (3a),(3b) and (4). The true values (true—) were calculated directly from the real data of model (3a), (3b) and (4). The values of $\langle\omega_{\mathrm{m}}\rangle$ and $\sum_{i=1}^{N}\Bigl{\langle}J_{Q_{\mathrm{b}}}^{(i)}\Bigr{\rangle}$ required for the calculation of the approximate values were computed using analytical solutions obtained from the quasilinear response relations (17), while $\langle P_{\mathrm{rel}}\rangle$ was calculated in three different ways, using Eqs. (24) (corresponding to approx1—), (26) (corresponding to approx2—) and (27) (corresponding to approx3—). The values of $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ and $\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$ contained in $\bigl{\langle}P_{\mathrm{rel}}\rangle$ were obtained from numerical experiments in the calculation of approximate values. (a): forward process, (b): backward process.
@@ III.5 Discussion
Before evaluating the approximation error by Eqs. (26) and (27), we first review the approximation procedure used to obtain the quasilinear response relations (17) (c.f. Appendix A). We approximated $\omega_{i}$ to a certain constant $\overline{\omega_{k}^{(i)}}$ each time the engine makes one rotation there since the variation of $\omega_{i}$ during one rotation of the phase $\theta_{i}$ is sufficiently small. Therefore, the approximation error that occurs when calculating the long-time average of the driving force portion of the engine or the heat flux is of the same order of magnitude as the time average of the minute approximation error that occurs for each rotation of the engine. On the other hand, Eqs. (26) and (27) are calculated via Eq. (28), which is obtained by approximating the value of $\omega_{i}$ to its long-time average $\langle\omega_{i}\rangle$, so the approximation accuracy is worse than that of relations (17). If the coupling strength is small enough, the value of $\omega_{i}$ at any given time fluctuates little from its long time average $\langle\omega_{i}\rangle$, so $\langle\omega_{i}\rangle$ is not much different from the time average of one rotation of $\theta_{i}$, but if the coupling strength is increased to some extent, the range of variation of $\omega_{i}$ on the long time scale becomes larger, and $\omega_{i}$ can deviate significantly from its long time average $\langle\omega_{i}\rangle$. Therefore, as confirmed by numerical experiments, the approximation errors of Eqs. (26) and (27) are larger than those of (24) in the region just before the synchronous transition occurs. After the synchronous transition occurs, the right-hand side of Eq.(28) becomes exactly zero and the approximation error due to the above approximation disappears.
@@
The above is a discussion of the approximation error resulting from the analysis conducted in subsection A from the computational point of view. To see the nature of the error from a physical point of view, let us look back again at the calculation of the entropy production rate (10)-(13) in subsection A. The intermediate formula (12) uses approximation (14), which is essentially the same as approximation (28). Since the power due to coupling is zero, the averaged entropy production rate for the total system should normally be expressed without using the long-time average of the coupling force. That is, the force conjugate to $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ should be $\frac{1}{N}\left(\sum_{k=1}^{N}T_{\mathrm{load}}^{(k)}\right)-T_{\mathrm{load}}^{(i)}$, not $\frac{1}{N}\left(\sum_{k=1}^{N}T_{\mathrm{load}}^{(k)}\right)-T_{\mathrm{load}}^{(i)}-\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$ exactly. However, this does not mean that the averaged entropy production rate is independent of the coupling strength. It is clear from the previous discussion that as the coupling strength of the total system is increased, the variance of $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ approaches 0, thereby reducing the entropy production of the total system. This means that if the force conjugate to $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ is chosen to be $\frac{1}{N}\left(\sum_{k=1}^{N}T_{\mathrm{load}}^{(k)}\right)-T_{\mathrm{load}}^{(i)}$, the coupling strength must be included in the response coefficients, which is the essence of irreversibility in the relative motion: the thermodynamic forces of load heterogeneity generates the conjugate fluxes of relative motion $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$, while the coupling strength is an internal parameter that regulates $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ caused by load heterogeneity. The stronger the coupling strength is, the greater the suppression of fluxes $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ due to load heterogeneity is. In this case, however, it is difficult to obtain the response coefficients analytically (in fact, the relations between the fluxes and forces are not even linear), so we dared to use the approximation (14) when calculating the averaged entropy production rate. Note that we included the term $\sum_{i=1}^{N}\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}\Bigl{\langle}\omega_{\rm d}^{(i)}\Bigr{\rangle}$ when computing $\bigl{\langle}P_{\rm rel}\bigr{\rangle}$ in Eq. (26), which corresponds to the above operation of putting $-\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$ in the force conjugate to $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$. Thus, the inaccuracy in the analysis of the entropy production and the power $\bigl{\langle}P_{\rm rel}\bigr{\rangle}$ using quasilinear response relations (17) in subsection A is due to the fact that the coupling strength, which is essentially an internal parameter that regulates the fluxes due to load heterogeneity, was put into the thermodynamic forces that produce irreversibility.
@@
Let us now discuss the approximation errors that arise in the analysis using the conceptual model constructed in subsection B. The analysis of the averaged power and thermal efficiency for this conceptual model does not go through any approximate calculations, so the results are completely accurate if we assume that the model itself is correct. As shown in subsection C, the differential equations and the heat flux equation for the conceptual models can be obtained by averaging the fast variables of the original coupled system models (3a), (3b) and (4). The time derivative of $\omega_{i}$ and the heat flux $J_{Q_{\mathrm{b}}}^{(i)}(\theta_{i},\omega_{i})$ were replaced by their averaged values during one cycle of $\theta_{i}$, and $\theta_{i}$ and $\theta_{i}-\theta_{j}$ were approximated as constants during one cycle of $\theta_{i}$. On the other hand, we used an approximate formula (47) when calculating the averaged entropy production rate for the isolated system D, which is also practically the same as the approximate formula (28). Therefore, the averaged entropy production rate (45) of system D is different from the averaged entropy production rate (46) determined by the given fluxes and the forces. However, if we note that in the original coupled engine model, the force conjugate to flux $\Bigl{\langle}\omega_{\rm d}^{(i)}\Bigr{\rangle}$, which should originally have been given by $\frac{1}{N}\left(\sum_{k=1}^{N}T_{\mathrm{load}}^{(k)}\right)-T_{\mathrm{load}}^{(i)}$, was replaced by $\frac{1}{N}\left(\sum_{k=1}^{N}T_{\mathrm{load}}^{(k)}\right)-T_{\mathrm{load}}^{(i)}-\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$, we find that Eq. (45), rather than Eq. (42), is appropriate as the averaged entropy production rate for system D. This imples that model D has been constructed as it should be. Unlike the quasilinear response relations (17), the conceptual model constructed in subsection B correctly reflects the nature of irreversibility in the relative motion. Thus, the conceptual model provides accurate analytical results for power and thermal efficiency without requiring the unknown, nonlinear relations that are tied between the real thermodynamic fluxes and forces.
@@
Although the synchronous transition exhibits the characteristics of a homoclinic bifurcation, we cannot conclude directly from the above analysis that the synchronous transition is caused by a homoclinic bifurcation. Therefore, we will analyze the process leading to the synchronous transition in more detail. If one-step increments of the coupling strength near $K_{\rm fd}$ are made finer, a phenomenon called the devil’s staircase [8], in which quasi-periodic and $m:n$ synchronous states $(m\neq n)$ appear alternately as the coupling strength is increased, can be observed. (Fig. 11(b)). The areas where $\langle\omega_{\rm d}\rangle$ is nearly flat (e.g. the areas that are enclosed by green dashed circles) exhibit $m:n$ synchronous states, which are limit cycles circling around the phase space $\mathbb{T}^{2}\times\mathbb{R}^{2}$, while the rest exhibit quasi-periodic states. Note that there exist many such nearly flat regions other than the areas enclosed by the green dashed lines, such as the areas between these enclosed areas. Like the $1:1$ synchronous states, a pair of $m:n$ synchronous states, one is stable and the other is unstable, are generated by a saddle-node bifurcation when the coupling strength $K$ is increased up to a certain threshold $K_{m:n}^{1}$. If we further increase the value of $K$, the flat part of the staircase collapses, suggesting that the above two $m:n$ synchronized states disappear due to a saddle-node bifurcation that occurs at another threshold $K_{m:n}^{2}$. The disappearance of the stable $m:n$ synchronous state causes the system in that state to either move to a quasi-periodic state or converge to another $m:n$ synchronous state. It should be noted that before the transition to a stable $m:n$ synchronous state occurs, that $m:n$ synchronous state has already been created by a saddle-node bifurcation. Accordingly, the flat parts of the devil’s staircase are narrower than the regions where the actual $m:n$ synchronous states can exist. Since Fig. 11(b) suggests that the $1:1$ synchronous transition is a discontinuous transition caused by the disappearance of a certain stable $m:n$ synchronous state at a saddle-node bifurcation, the synchronous transition is not strictly a homoclinic bifurcation. To be precise, the process leading to the $1:1$ synchronous transition is the one in which the transitions of an asynchronous state (quasi-periodic or an $m:n$ synchronous state) to another $m:n$ synchronous state and the collapse of $m:n$ synchronous states due to saddle node bifurcations occur alternately while reserving the feature of a homoclinic bifurcation in that the asynchronous states gradually approach the saddle limit cycle corresponding to the unstable $1:1$ synchronous state. However, if $\Delta\omega_{\rm n}$ is sufficiently small compared to $\omega_{\rm n}^{(i)}$, the $m:n$ synchronous regions turn to be sufficiently narrow compared to the range of the variation of the coupling strength $K$ up to $K_{\rm fd}$, so on a macroscopic scale, the synchronous transition can be considered to be caused by a homoclinic bifurcation approximately. If the load torques are adjusted so that $\Delta\omega_{\rm n}$ is made larger, the nearly flat $m:n$ synchronous regions in the devil’s staircase become wider, and the coarse-grained homoclinic bifurcation structure in the synchronous transition breaks down. Figure 11(c) shows the dependence relation between $\langle\omega_{\rm d}\rangle$ and $K$ with the value of $\Delta\omega_{\mathrm{n}}$ slightly larger than that in Fig. 11(a). The areas corresponding to $m:n$ synchronous states are wider than those observed in Fig. 11(a), and it can be seen even on a macroscopic scale that the synchronous transition occurs discontinuously. Furthermore, the $m:n$ synchronous regions observed in the devil’s staircase tend to become wider as the synchronous transition point is approached. This dependence of the width of the $m:n$ synchronous regions on $\Delta\omega_{\mathrm{n}}$ and $K$ can be explained by the distorted V-shaped structure of Arnold’s tongues [8].
@@
Next, we consider the dynamics of the conceptual model. The equations describing the relative motion, neglecting higher-order terms, are given by

### v2

@@ III.3 Justification of the conceptual model
In this section, we show that the differential equations describing the rotational motion of the conceptual model constructed in subsection B can be obtained by eliminating fast oscillating variables in those of the original coupled Stirling engine model. Since conditions (5) and (7) ensure that the time scales of $\omega_{i}$ and $\theta_{i}-\theta_{j}$ are sufficiently large compared to that of $\theta_{i}$, i.e., $\omega_{i}$ and $\theta_{i}-\theta_{j}$ change only slightly during which $\theta_{i}$ varies by $2\pi$, if we replace the time derivative of $\omega_{i}$ by its time average during a one-round change in $\theta_{i}$, $\omega_{i}$ and $\theta_{i}-\theta_{j}$ can be regarded as constants so that we can just replace the driving force portion of the engine by its time average over one cycle of $\theta_{i}$ in Eq. (3b):
@@
The higher-order terms produced by the averaging approximation of the heat flux can be ignored given that conditions (5) and (7) are satisfied.
The above discussion confirms the validity of the conceptual model constructed in subsection B, but there are still several issues to be addressed. One is that the results of the analysis in subsection B using the conceptual model do not agree with the results of the analysis in subsection A using the quasilinear response relations (17): the averaged power $\Bigl{\langle}P_{\mathrm{load}}^{\mathrm{D}}\Bigr{\rangle}$ applied by the flywheels ${\mathrm{D}}_{\mathrm{F}W}^{(1)},{\mathrm{D}}_{\mathrm{F}W}^{(2)},\cdots,{\mathrm{D}}_{\mathrm{F}W}^{(N)}$ to the loads (or the averaged power $-\Bigl{\langle}P_{\mathrm{fric}}^{\mathrm{D}}\Bigr{\rangle}$ due to the viscous friction) is given by Eq. (43) with the sign reversed:
@@
which corresponds to the averaged brake power $\langle P_{\mathrm{rel}}\rangle$ due to the relative motion of the weakly-coupled Stirling engines, but this result is different from the approximation $\langle P_{\mathrm{rel}}\rangle\approx-L_{11}^{-1}\left(\sum_{i=1}^{N}\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}^{2}\right)$ given by Eq. (27). The reason why the analysis based on the conceptual model has different results from the analysis using the quasilinear response relations (17) will be discussed in subsection E. Another issue is the difference in energy dissipation due to viscous friction between the conceptual model and the original coupled engine model: the averaged energy dissipation rate due to viscous friction of the coupled Stirling engine is given by $\sum_{i=1}^{N}\Gamma\omega_{i}^{2}$, whereas in the conceptual model it is given by $\left(\Gamma+\frac{\sigma^{2}}{G}\Biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\Biggr{\rangle}_{\theta}\right)\left[N\omega_{\mathrm{m}}^{2}+\sum_{i=1}^{N}\left(\omega_{\mathrm{d}}^{(i)}\right)^{2}\right]=\left(\Gamma+\frac{\sigma^{2}}{G}\Biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\Biggr{\rangle}_{\theta}\right)\left(\sum_{i=1}^{N}\omega_{i}^{2}\right)$. This is due to the fact that the averaged energy dissipation rate $\frac{\sigma^{2}}{G}\Biggl{\langle}\frac{\sin^{2}\theta}{V^{2}(\theta)}\Biggr{\rangle}_{\theta}\left(\sum_{i=1}^{N}\omega_{i}^{2}\right)$ in the conceptual model comes from the heat leakage of the coupled Stirling engines (See Appendix B for details).
@@ III.4 Numerical experiments
Figure 5: Dependence of the coupling strength $K$ on the averaged brake power due to relative motion $\langle P_{\mathrm{rel}}\rangle$, calculated for coupled system models (3a)-(3b). The true values (true—) were calculated from model (3a)-(3b), while the approximate values (approx1—, approx2— and approx3—) were calculated from Eqs. (24), (26), and (27), respectively. The values of $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ and $\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$ contained in $\bigl{\langle}P_{\mathrm{rel}}\rangle$ were obtained from numerical experiments in the calculation of approximate values. (a): forward process, (b): backward process.
@@
Figure 6: Dependence of the coupling strength $K$ on the averaged brake power $\langle P_{\mathrm{load}}\rangle$, calculated for coupled system models (3a)-(3b). The true values (true—) were calculated from model (3a)-(3b). The values of $\langle\omega_{\mathrm{m}}\rangle$ required for the calculation of the approximate values were computed using analytical solutions obtained from the quasilinear response relations (17), while $\langle P_{\mathrm{rel}}\rangle$ was calculated in three different ways, using Eqs. (24) (corresponding to approx1—), (26) (corresponding to approx2—) and (27) (corresponding to approx3—). The values of $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ and $\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$ contained in $\bigl{\langle}P_{\mathrm{rel}}\rangle$ were obtained from numerical experiments in the calculation of approximate values. (a): forward process, (b): backward process.
@@
Figure 7: Dependence of the coupling strength $K$ on the thermal efficiency $\eta$, calculated for coupled system models (3a),(3b) and (4). The true values (true—) were calculated from model (3a), (3b) and (4). The values of $\langle\omega_{\mathrm{m}}\rangle$ and $\sum_{i=1}^{N}\Bigl{\langle}J_{Q_{\mathrm{b}}}^{(i)}\Bigr{\rangle}$ required for the calculation of the approximate values were computed using analytical solutions obtained from the quasilinear response relations (17), while $\langle P_{\mathrm{rel}}\rangle$ was calculated in three different ways, using Eqs. (24) (corresponding to approx1—), (26) (corresponding to approx2—) and (27) (corresponding to approx3—). The values of $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ and $\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$ contained in $\bigl{\langle}P_{\mathrm{rel}}\rangle$ were obtained from numerical experiments in the calculation of approximate values. (a): forward process, (b): backward process.
@@ III.5 Discussion
Before evaluating the approximation error by Eqs. (26) and (27), we first review the approximation procedure used to obtain the quasilinear response relations (17) (c.f. Appendix A). We approximated $\omega_{i}$ to a certain constant $\overline{\omega_{k}^{(i)}}$ each time the engine makes one rotation there since the variation of $\omega_{i}$ during one rotation of the phase $\theta_{i}$ is sufficiently small. Therefore, the approximation error that occurs when calculating the long-time average of the driving force portion of the engine or the heat flux is of the same order of magnitude as the time average of the minute approximation error that occurs for each rotation of the engine. In this sense, the quasilinear response relations (17) are mathematically valid relations. On the other hand, Eqs. (26) and (27) are calculated via Eq. (28), which is obtained by approximating the value of $\omega_{i}$ to its long-time average $\langle\omega_{i}\rangle$, so the approximation accuracy is worse than that of relations (17). If the coupling strength is small enough, the value of $\omega_{i}$ at any given time fluctuates little from its long time average $\langle\omega_{i}\rangle$, so $\langle\omega_{i}\rangle$ is not much different from the time average of one rotation of $\theta_{i}$, but if the coupling strength is increased to some extent, the range of variation of $\omega_{i}$ on the long time scale becomes larger, and $\omega_{i}$ can deviate significantly from its long time average $\langle\omega_{i}\rangle$. Therefore, as confirmed by numerical experiments, the approximation errors of Eqs. (26) and (27) are larger than those of (24) in the region just before the synchronous transition occurs. After the synchronous transition occurs, the right-hand side of Eq.(28) becomes exactly zero and the approximation error due to the above approximation disappears.
@@
The above is a discussion of the approximation error of the analysis based on the quasilinear response relations (17) from a mathematical point of view. To give a discussion from the point of view of nonequilibrium thermodynamics, let us look back again at the calculation of the entropy production rate (10)-(13) in subsection A. The intermediate formula (12) uses approximation (14), which is essentially the same as approximation (28). Since the averaged power exerted by the coupling forces between engines on the total coupled system is zero, the averaged entropy production rate for the total system should normally be expressed without using the long-time average of the coupling force. That is, the real thermodynamic force conjugate to $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ should be the one expressing the load heterogeneity, $\frac{1}{N}\left(\sum_{k=1}^{N}T_{\mathrm{load}}^{(k)}\right)-T_{\mathrm{load}}^{(i)}$, not $\frac{1}{N}\left(\sum_{k=1}^{N}T_{\mathrm{load}}^{(k)}\right)-T_{\mathrm{load}}^{(i)}-\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$ exactly, and the coupling strength should be included in the response coefficients. This implies that the coupling strength is an internal parameter of the coupled system that regulates the fluxes $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ caused by the load heterogeneity. The stronger the coupling strength is, the greater the suppression of fluxes $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ due to the load heterogeneity is. In this case, however, it is difficult to obtain the response coefficients analytically (in fact, the relations between the fluxes and forces are not even linear, and show hysteresis structures that originate from the dependence on the state variables of the coupled system), so we dared to use the approximation (14) and included the averaged coupling force $-\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$ in the thermodynamic forces conjugate to the fluxes $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$ when calculating the averaged entropy production rate. On the other hand, we included the term $\sum_{i=1}^{N}\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}\Bigl{\langle}\omega_{\rm d}^{(i)}\Bigr{\rangle}$ when computing $\bigl{\langle}P_{\rm rel}\bigr{\rangle}$ in Eq. (26), which corresponds to the above operation of putting $-\sum_{j\in\mathcal{N}_{i}}K_{ij}\Bigl{\langle}\sin\left(\theta_{i}-\theta_{j}\right)\Bigr{\rangle}$ in the force conjugate to $\Bigl{\langle}\omega_{\mathrm{d}}^{(i)}\Bigr{\rangle}$. Thus, the inaccuracy in the analysis of the averaged brake power and the thermal efficiency using quasilinear response relations (17) in subsection A is due to the fact that the coupling strength, which is essentially an internal parameter that regulates the fluxes due to load heterogeneity, was put into the thermodynamic forces that produce irreversibility. In other words, the discrepancy between the analytical results and the true values is due to the fact that the quasilinear response relations (17) do not preserve the thermodynamic irreversibility of the original coupled system.
@@
Let us now discuss the approximation errors that arise in the analysis using the conceptual model constructed in subsection B. The analysis of the averaged power and thermal efficiency for this conceptual model does not go through any approximate calculations, so the results are completely accurate if we assume that the model itself is valid. The conceptual model was constructed to satisfy the quasilinear response relations (17), but since the relations include the terms of the averaged coupling forces, which depend on the dynamics of the original coupled engines, we cannot conclude directly that the conceptual model satisfies the quasilinear response relations (17) unless the terms of the averaged coupling forces in the conceptual model are in close agreement with those in the original coupled engine model. As shown in subsection C, the terms of the averaged coupling forces in the conceptual model and the original coupled model are nearly identical given that conditions (5), (7) and (8) are satisfied, so the conceptual model satisfies the quasilinear response relations (17) with good approximation accuracy. On the other hand, since the quasilinear response relations (17) do not preserve the thermodynamic irreversibility of the original coupled system, a concern arises that the conceptual model constructed from the quasilinear response relations may not preserve the thermodynamic irreversibility as well. However, the averaged entropy production rate of the constructed isolated system D is given by Eq. (44), which is not consistent with that determined from the quasilinear response relations (30). The approximate formula (46), which is practically equivalent to the approximate formula (28), was used in the calculation of the averaged entropy production rate for the isolated system D to ensure that system D satisfies the quasilinear response relations (30). It can be seen from the averaged entropy production rate (44) for the isolated system D that the thermodynamic fluxes and forces of the conceptual model are in perfect agreement with the real thermodynamic fluxes and forces of the original coupled system if the averaging error is ignored, and thus, unlike the quasilinear response relations (17), the conceptual model constructed in subsection B correctly reflects the nature of the thermodynamic irreversibility of the original coupled system. This is the reason why the analytical results using the quasilinear response relations (17) differ from those using the conceptual model. Therefore, the conceptual model provides accurate analytical results for the averaged power and thermal efficiency without requiring the unknown response relations between the real thermodynamic fluxes and forces.
@@
Although the synchronous transition exhibits the characteristics of a homoclinic bifurcation, we cannot conclude directly from the above analysis that the synchronous transition is caused by a homoclinic bifurcation. Therefore, we will analyze the process leading to the synchronous transition in more detail. If one-step increments of the coupling strength near $K_{\rm fd}$ are made finer, a phenomenon called the devil’s staircase [8], in which quasi-periodic and $m:n$ synchronous states $(m\neq n)$ appear alternately as the coupling strength is increased, can be observed. (Fig. 11(b)). The areas where $\langle\omega_{\rm d}\rangle$ is nearly flat (e.g. the areas that are enclosed by green dashed circles) exhibit $m:n$ synchronous states, which are limit cycles circling around the phase space $\mathbb{T}^{2}\times\mathbb{R}^{2}$, while the rest exhibit quasi-periodic states. Note that there exist many such nearly flat regions other than the areas enclosed by the green dashed lines, such as the areas between these enclosed areas. Like the $1:1$ synchronous states, a pair of $m:n$ synchronous states, one is stable and the other is unstable, are generated by a saddle-node bifurcation when the coupling strength $K$ is increased up to a certain threshold $K_{m:n}^{1}$. If we further increase the value of $K$, the flat part of the staircase collapses, suggesting that the above two $m:n$ synchronized states disappear due to a saddle-node bifurcation that occurs at another threshold $K_{m:n}^{2}$. The disappearance of the stable $m:n$ synchronous state causes the system in that state to either move to a quasi-periodic state or converge to another $m:n$ synchronous state. It should be noted that before the transition to a stable $m:n$ synchronous state occurs, that $m:n$ synchronous state has already been created by a saddle-node bifurcation. Accordingly, the flat parts of the devil’s staircase are narrower than the regions where the actual $m:n$ synchronous states can exist. Since Fig. 11(b) suggests that the $1:1$ synchronous transition is a discontinuous transition caused by the disappearance of a certain stable $m:n$ synchronous state at a saddle-node bifurcation, the synchronous transition is not strictly a homoclinic bifurcation. To be precise, the process leading to the $1:1$ synchronous transition is the one in which the transitions of an asynchronous state (quasi-periodic or an $m:n$ synchronous state) to another $m:n$ synchronous state and the collapse of $m:n$ synchronous states due to saddle node bifurcations occur alternately while reserving the feature of a homoclinic bifurcation in that the asynchronous states gradually approach the saddle limit cycle corresponding to the unstable $1:1$ synchronous state. However, if $\Delta\omega_{\rm n}$ is sufficiently small compared to $\omega_{\rm n}^{(i)}$, the $m:n$ synchronous regions turn to be sufficiently narrow compared to the range of the variation of the coupling strength $K$ up to $K_{\rm fd}$, so on a macroscopic scale, the synchronous transition can be considered to be caused by a homoclinic bifurcation approximately. If the load torques are adjusted so that $\Delta\omega_{\rm n}$ is made larger, the nearly flat $m:n$ synchronous regions in the devil’s staircase become wider, and the coarse-grained homoclinic bifurcation structure in the synchronous transition breaks down. Figure 11(c) shows the dependence relation between $\langle\omega_{\rm d}\rangle$ and $K$ with the value of $\Delta\omega_{\mathrm{n}}$ slightly larger than that in Fig. 11(a). The areas corresponding to $m:n$ synchronous states are wider than those observed in Fig. 11(a), and it can be seen even on a macroscopic scale that the synchronous transition occurs discontinuously. Furthermore, the $m:n$ synchronous regions observed in the devil’s staircase tend to become wider as the synchronous transition point is approached in the forward process. This dependence of the width of the $m:n$ synchronous regions on $\Delta\omega_{\mathrm{n}}$ and $K$ can be explained by the distorted V-shaped structure of Arnold’s tongues [8].
@@
Next, we consider the differential equations of the conceptual model (i.e., the averaged differential equations w.r.t. Eqs. (3a) and (3b)). The equations describing the rotational motion of subsystem D are given by